# HHardware specifications

### v1

@@ 5 Conclusion
Future work will extend this study in several directions. This framework could readily be used to estimate the energy consumption of photonic computation architectures, such as [41]. More consideration must be given to the cost associated to classical post-processing, and in particular to the trade-off between the energy spent in post-processing and the level of security. Since said task is more involved for CV protocols, this might give more insight into which method is more efficient. Furthermore, it is worth considering the authentication cost of the classical channels used by the parties. This is done using pre-shared keys [42] or post-quantum cryptography [43], and is a requirement to avoid man-in-the-middle attacks. The costs of these tasks are difficult to estimate because there is no standard method to perform them yet. Finally, the list of hardware available in the accompanying software can be expanded, including for example new components such as quantum memories or solid-state sources. This will contribute to rigorous benchmarking and optimization of the energy consumption of large scale quantum networks with heterogeneous hardware for each nodes and hybrid fiber/free-space links.
@@ Acknowledgements
The authors thank Paul Hilaire, Simon Neves and Verena Yacoub for fruitful discussions and inputs, as well as Eleni Diamanti for her feedback and guidance.
@@ Appendix A Table of components
Theoretical values for $E_{0}$ are obtained by multiplying the initialization time by their power. Measured $E_{0}$ is obtained either by multiplying the initialization time by the measured power, or fully measured by following the power consumption in real time and integrating (marked by an asterisk *).
For the measured values in Table 3 (column 5 and 7), the measurements were done in one of the three following ways, depending on its power connection type:
@@
  • If the device was plugged to a standard lab power supply, then the measurement was done by applying the required voltage, and recording the consumed current, and multiplying the two values to get the power.
@@
When possible the power consumption was measured when the device was being used. The measurements for the lasers were done while emitting their optical beam; for the single photon detectors, while detecting a flow of single photons; for the computers, while they were running the control software of the CV-QKD experiment; for the waveplates, while they were rotating; for the AM modulator, while applying a pulsed signal (for the signal generator) and while locking (for the modulator bias controller); for the oven, while stabilizing a temperature of 25C; for the IQ modulator, while locking the modulator; for the powermeter, while receiving optical power; for for the optical switch, while applying switching commands; for the ADC and DAC, while running the CV-QKD experiment (and hence emitting and receiving signals).
Additionally for some equipment with a long initialization time (such as lasers or single photon detectors), the power consumption was recorded during the initialization.
@@ B.2.1 Weak coherent pulse sources
Weak coherent pulse sources are a practical and efficient method to obtain single-photons probabilistically. The idea consists of attenuating a laser pulse until the probability that a light packet contains more than one photon is low enough that it can be neglected or effectively bounded in security proofs. From a hardware point of view, this requires a laser, passive attenuation and, usually, an amplitude modulator.
@@ B.3 Manipulation
$\left|\phi^{+}\right\rangle_{A}\otimes\left|\phi^{+}\right\rangle_{B}$ $=\frac{1}{\sqrt{2}}(\left|0000\right\rangle_{A_{1}A_{2}B_{1}B_{2}}+\left|1100\right\rangle_{A_{1}A_{2}B_{1}B_{2}}+\left|0011\right\rangle_{A_{1}A_{2}B_{1}B_{2}}+\left|1111\right\rangle_{A_{1}A_{2}B_{1}B_{2}})$
$\rightarrow\frac{1}{\sqrt{2}}(\left|0000\right\rangle_{A_{1}xyB_{2}}+\left|1100\right\rangle_{A_{1}yyB_{2}}+\left|0011\right\rangle_{A_{1}xxB_{2}}+\left|1111\right\rangle_{A_{1}xyB_{2}}).$
@@
Figure 16: Schematics of setups to create and share 4-qubit GHZ states encoded a) in polarization and b) in time.
@@ B.4.2 Coherent detectors
For CV-QKD with homodyne detection, a Motorized Polarization Controller (MPC) and a Switch (S), used respectively to compensate long-distance polarization dispersion and to calibrate the noise levels, receive the light signal and mix it with a Local Oscillator (LO) in a beam splitter. The result is sent to the BHD device. It also requires a phase modulator on the LO path in order to select the measured quadrature (and perform a sifted protocol, as in BB84 for instance). An Analog-to-Digital Converter (ADC) allows the acquisition of data. For heterodyne detection, the general concept is the same, except that the signal light is divided in two. Each of these outputs is separately mixed with the local oscillator, one of the two being dephased by $\pi/2$, before being sent to a BHD each, allowing for simultaneous measurement of both quadratures. The two possible detection schemes are presented in Figures 17(a) and 17(b).
@@
$\begin{split}D_{\rm CV-QKD,hom,1P}(t)&=t(P_{\rm ADC}+P_{\rm laser}+P_{\rm BHD}+P_{\rm PC}+P_{\rm PM}),\\ D_{\rm CV-QKD,het,1P}(t)&=t(P_{\rm ADC}+P_{\rm laser}+2P_{\rm BHD}+P_{\rm PC}),\\ D_{\rm CV-QKD,hom,2P}(t)&=t(P_{\rm ADC}+P_{\rm laser}+2P_{\rm BHD}+P_{\rm PC}+P_{\rm PM}),\\ D_{\rm CV-QKD,het,2P}(t)&=t(P_{\rm ADC}+P_{\rm laser}+4P_{\rm BHD}+P_{\rm PC}),\\ \end{split}$ (28)
@@ C.1.1 BB84
$E_{\textrm{BB84, polar/time}}(t)=S_{\textrm{weak, polar/time}}(t)+2\;M_{\rm polar/time}(t)+D_{\textrm{APD/SNSPD}}(t)+C_{\textrm{BB84}}(t).$ (31)
@@ C.1.2 E91
$E_{\textrm{E91, time/polar}}(t)=S_{\textrm{SPDC, time/polar}}(t)+2\;M_{\rm time/polar}(t)+2\;D_{\textrm{APD/SNSPD}}(t)+C_{\textrm{E91}}(t).$ (33)
@@ C.1.3 MDI-QKD
$M_{BSM,polarization}(t)$ $=0$
$M_{BSM,time}(t)$ $=2tP_{interferometry}$ (35)
@@
$E_{\textrm{MDI}}(t)=2\;S_{\textrm{weak,polar/time}}(t)+M_{\rm BSM,polar/time}(t)+D_{\textrm{APD/SNSPD}}(t)+C_{\textrm{MDI}}(t).$ (36)
@@ C.2.2 Computation of the secret key rate of CV-QKD protocols
$I_{AB,hom}=\frac{1}{2}\log_{2}(1+\text{SNR})=\frac{1}{2}\log_{2}\left(1+\frac{p_{\rm det}TV_{A}}{1+V_{el}+p_{\rm det}T\xi}\right),$ (40)
@@
$I_{AB,het}=\log_{2}(1+\text{SNR})=\log_{2}\left(1+\frac{p_{\rm det}TV_{A}}{2+2V_{el}+p_{\rm det}T\xi}\right),$ (41)
@@
$\begin{split}\lambda^{2}_{3,4}&=\frac{1}{2}\left[C\pm\sqrt{C^{2}-4D}\right],\\ C_{hom}&=\frac{A\chi_{hom}+V\sqrt{B}+T(V+\chi_{line})}{T(V+\chi_{tot,hom})},\\ D_{hom}&=\sqrt{B}\frac{V+\sqrt{B}\chi_{hom}}{T(V+\chi_{tot,hom})},\\ C_{het}&=\frac{1}{(V(T+\chi_{tot,het}))^{2}}\left[A\chi_{het}^{2}+B+1+2\chi_{het}(V\sqrt{B}+T(V+\chi_{line}))+2T(V^{2}-1)\right],\\ D_{het}&=\left(\frac{V+\sqrt{B}\chi_{het}}{T(V+\chi_{tot,het})}\right)^{2}.\end{split}$ (45)
@@
$\begin{split}V&=V_{A}+1,\\ W&=1+p_{\rm det}TV_{A}+p_{\rm det}T\xi+V_{el},\\ Z&=\sqrt{T}\left(2\alpha^{2}e^{-\alpha^{2}}\sum_{k=0}^{M-1}\frac{\nu_{k}^{3/2}}{\nu_{k+1}^{1/2}}-\sqrt{2\xi\alpha^{2}}\sqrt{e^{-\alpha^{2}}\sum_{j=0}^{M-1}\frac{\nu_{j}^{2}}{\nu_{j+1}}-e^{-2\alpha^{2}}\left(\sum_{j=0}^{M-1}\frac{\nu_{j}^{3/2}}{\nu_{j+1}^{1/2}}\right)^{2}}\right),\end{split}$ (46)
@@ C.2.3 Computation of the secret key rate of the CV-CKA protocol
$\Delta=\mathrm{diag}\{y-(n-1)z^{2}/(nx),y-z^{2}/(nx)\},$ $\Theta=\mathrm{diag}\{z^{2}/(nx),-z^{2}/(nx)\},$
@@ Appendix D Time-bin encoding
The same result holds for the other DV protocols considered in this study: using time-bin encoding results in a slight increase in the energy consumption. Due to the lack of major differences, time encoded plots for QKD and CKA were excluded from Figures 1, 2 and 8 for readability.

### v2

@@ V Conclusion
Future work will extend this study in several directions. This framework could readily be used to estimate the energy consumption of photonic computation architectures, such as [44]. More consideration must be given to the cost associated with classical post-processing, and in particular to the trade-off between the energy spent in post-processing and the level of security. Since said task is more involved in CV protocols, this might give more insight into which method is more efficient. Furthermore, it is worth considering the authentication cost of the classical channels used by the parties. This is done using pre-shared keys [45] or post-quantum cryptography [46], and is a requirement to avoid man-in-the-middle attacks. The cost of these tasks is difficult to estimate because there is no standard method to perform them yet. Finally, the list of hardware available in the accompanying software can be expanded, including for example new components such as quantum memories or solid-state sources. This will contribute to rigorous benchmarking and optimization of the energy consumption of large scale quantum networks with heterogeneous hardware for each nodes and hybrid fiber/free-space links.
@@ Acknowledgements
The authors thank Paul Hilaire, Simon Neves and Verena Yacoub for fruitful discussions and inputs, as well as Eleni Diamanti and Alexia Auffèves for their feedback and guidance.
@@ Appendix A Table of components
Laser $\lambda$ (nm) $E_{0}$ (kJ) Meas. (kJ) $P$ (W) Meas. (W) Ref.
[…]
Detector $\lambda$ (nm) $E_{0}$ (kJ) Meas. (kJ) $P$ (W) Meas. (W) Ref.
[…]
Component $E_{0}$ (kJ) Meas. (kJ) $P$ (W) Meas. (W) Ref.
[…]
@@
Theoretical values for $E_{0}$ are obtained by multiplying the initialization time by their power. Measured $E_{0}$ is obtained either by multiplying the initialization time by the measured power, or fully measured by following the power consumption in real-time and integrating (marked by an asterisk *).
For the measured values in Table 3 (columns 5 and 7), the measurements were done in one of the three following ways, depending on its power connection type:
@@
  • If the device was plugged into a standard lab power supply, then the measurement was done by applying the required voltage, and recording the consumed current, and multiplying the two values to get the power.
@@
When possible the power consumption was measured while the device was being used. The measurements for the lasers were done while emitting their optical beam; for the single-photon detectors, while detecting a flow of single photons; for the computers, while they were running the control software of the CV-QKD experiment; for the waveplates, while they were rotating; for the AM modulator, while applying a pulsed signal (for the signal generator) and while locking (for the modulator bias controller); for the oven, while stabilizing a temperature of 25C; for the IQ modulator, while locking the modulator; for the powermeter, while receiving optical power; for the optical switch, while applying switching commands; for the ADC and DAC, while running the CV-QKD experiment (and hence emitting and receiving signals).
Additionally, for some equipment with a long initialization time (such as lasers or single photon detectors), the power consumption was recorded during the initialization.
@@ B.2.1 Weak coherent pulse sources
Weak coherent pulse sources are a practical and efficient method to obtain single photons probabilistically. The idea consists of attenuating a laser pulse until the probability that a light packet contains more than one photon is low enough that it can be neglected or effectively bounded in security proofs. From a hardware point of view, this requires a laser, passive attenuation and, usually, an amplitude modulator.
@@ B.3 Manipulation
$\left|\phi^{+}\right\rangle_{A}\otimes\left|\phi^{+}\right\rangle_{B}$ $=\frac{1}{\sqrt{2}}\left(\left|0000\right\rangle_{A_{1}A_{2}B_{1}B_{2}}+\left|1100\right\rangle_{A_{1}A_{2}B_{1}B_{2}}\right.$
$\hskip 17.07182pt+\left.\left|0011\right\rangle_{A_{1}A_{2}B_{1}B_{2}}+\left|1111\right\rangle_{A_{1}A_{2}B_{1}B_{2}}\right)$
$\rightarrow\frac{1}{\sqrt{2}}\left(\left|0000\right\rangle_{A_{1}xyB_{2}}+\left|1100\right\rangle_{A_{1}yyB_{2}}\right.$
$\hskip 17.07182pt\left.+\left|0011\right\rangle_{A_{1}xxB_{2}}+\left|1111\right\rangle_{A_{1}xyB_{2}}\right).$
@@
(a) GHZ in polarization
(b) GHZ in time
Figure 17: Schematics of setups to create and share 4-qubit GHZ states encoded (a) in polarization and (b) in time.
@@ B.4.2 Coherent detectors
For CV-QKD with homodyne detection, a Motorized Polarization Controller (MPC) and a Switch (S), used respectively to compensate long-distance polarization dispersion and to calibrate the noise levels, receive the light signal and mix it with a Local Oscillator (LO) in a beam splitter. The result is sent to the BHD device. It also requires a phase modulator on the LO path in order to select the measured quadrature (and perform a sifted protocol, as in BB84 for instance). An Analog-to-Digital Converter (ADC) allows the acquisition of data. For heterodyne detection, the general concept is the same, except that the signal light is divided in two. Each of these outputs is separately mixed with the local oscillator, one of the two being dephased by $\pi/2$, before being sent to a BHD each, allowing for simultaneous measurement of both quadratures. The two possible detection schemes are presented in Figures 18(a) and 18(b).
@@
$\begin{split}D_{\rm CV,hom,1P}(t)&=t(P_{\rm ADC}+P_{\rm laser}+P_{\rm BHD}+P_{\rm PC}+P_{\rm PM}),\\ D_{\rm CV,het,1P}(t)&=t(P_{\rm ADC}+P_{\rm laser}+2P_{\rm BHD}+P_{\rm PC}),\\ D_{\rm CV,hom,2P}(t)&=t(P_{\rm ADC}+P_{\rm laser}+2P_{\rm BHD}+P_{\rm PC}+P_{\rm PM}),\\ D_{\rm CV,het,2P}(t)&=t(P_{\rm ADC}+P_{\rm laser}+4P_{\rm BHD}+P_{\rm PC}),\\ \end{split}$ (27)
@@ C.1.1 BB84
$\begin{split}E_{\textrm{BB84, polar/time}}(t)&=S_{\textrm{weak, polar/time}}(t)\\ &\hskip 11.38092pt+2\;M_{\rm polar/time}(t)\\ &\hskip 14.22636pt+D_{\textrm{APD/SNSPD}}(t)\\ &\hskip 17.07182pt+C_{\textrm{BB84}}(t).\end{split}$ (30)
@@ C.1.2 E91
$\begin{split}E_{\textrm{E91, time/polar}}(t)&=S_{\textrm{SPDC, time/polar}}(t)\\ &\hskip 11.38092pt+2\;M_{\rm time/polar}(t)\\ &\hskip 14.22636pt+2\;D_{\textrm{APD/SNSPD}}(t)\\ &\hskip 17.07182pt+C_{\textrm{E91}}(t).\end{split}$ (32)
@@ C.1.3 MDI-QKD
$M_{\rm BSM,polarization}(t)$ $=0$
$M_{\rm BSM,time}(t)$ $=2tP_{\rm interferometry}$ (34)
@@
$\begin{split}E_{\textrm{MDI}}(t)&=2\;S_{\textrm{weak,polar/time}}(t)\\ &\hskip 11.38092pt+M_{\rm BSM,polar/time}(t)\\ &\hskip 14.22636pt+D_{\textrm{APD/SNSPD}}(t)\\ &\hskip 17.07182pt+C_{\textrm{MDI}}(t).\end{split}$ (35)
@@ C.2.2 Computation of the secret key rate of CV-QKD protocols
$\begin{split}I_{AB,hom}={}&\frac{1}{2}\log_{2}(1+\text{SNR})\\ ={}&\frac{1}{2}\log_{2}\left(1+\frac{p_{\rm det}TV_{A}}{1+V_{el}+p_{\rm det}T\xi}\right),\end{split}$ (39)
@@
$\begin{split}I_{AB,het}={}&\log_{2}(1+\text{SNR})\\ ={}&\log_{2}\left(1+\frac{p_{\rm det}TV_{A}}{2+2V_{el}+p_{\rm det}T\xi}\right),\end{split}$ (40)
@@
$\begin{split}\lambda^{2}_{3,4}&=\frac{1}{2}\left[C\pm\sqrt{C^{2}-4D}\right],\\ C_{hom}&=\frac{A\chi_{hom}+V\sqrt{B}+T(V+\chi_{line})}{T(V+\chi_{tot,hom})},\\ D_{hom}&=\sqrt{B}\frac{V+\sqrt{B}\chi_{hom}}{T(V+\chi_{tot,hom})},\\ C_{het}&=\frac{1}{(V(T+\chi_{tot,het}))^{2}}\left[A\chi_{het}^{2}+B+1\right.\\ &\left.+2\chi_{het}(V\sqrt{B}+T(V+\chi_{line}))+2T(V^{2}-1)\right],\\ D_{het}&=\left(\frac{V+\sqrt{B}\chi_{het}}{T(V+\chi_{tot,het})}\right)^{2}.\end{split}$ (44)
@@
$\begin{split}V&=V_{A}+1,\\ W&=1+p_{\rm det}TV_{A}+p_{\rm det}T\xi+V_{el},\\ Z&=\sqrt{T}\left(2\alpha^{2}e^{-\alpha^{2}}\sum_{k=0}^{M-1}\frac{\nu_{k}^{3/2}}{\nu_{k+1}^{1/2}}\right.\\ &-\left.\sqrt{2\xi\alpha^{2}}\sqrt{e^{-\alpha^{2}}\sum_{j=0}^{M-1}\frac{\nu_{j}^{2}}{\nu_{j+1}}-e^{-2\alpha^{2}}\left(\sum_{j=0}^{M-1}\frac{\nu_{j}^{3/2}}{\nu_{j+1}^{1/2}}\right)^{2}}\right),\end{split}$ (45)
@@ C.2.3 Computation of the secret key rate of the CV-CKA protocol
$\Delta$ $=\mathrm{diag}\{y-(n-1)z^{2}/(nx),y-z^{2}/(nx)\},$
$\Theta$ $=\mathrm{diag}\{z^{2}/(nx),-z^{2}/(nx)\},$
@@ Appendix D Time-bin encoding
The same result holds for the other DV protocols considered in this study: using time-bin encoding results in a slight increase in the energy consumption. Due to the lack of major differences, time encoded plots for QKD and CKA were excluded from Figures 1, 2 and 9 for readability.